OpenStudy (anonymous):

A particular photon has a wavelength of approximately 2x10^-7 m. What is the energy, in joules, associated with this photon? a. 4 x 10^-32 b.1 x 10^-18 c. 1 x 10^18 d. More information is needed. e.1.5 x 10^15

OpenStudy (danjs):

Recall: The energy is the same as planks constant times the speed of light, divided by the wavelength \[Energy ~of~ Photon = \frac{ h*c }{ \lambda }\] \[h = 6.626*10^{-34} ~J*s\] c = 2.998 x 10^8 m/s so h*c = 1.99 x 10^-25 J*m

OpenStudy (anonymous):

@DanJS Thanks for the reply. In that case, I would do (1.99 x 10^-25 J*m)/( 2 x 10^-7). However, the answer would be 9.95 x 10^-19. That's where I'm confused, as that isn't one of the answer choices.

OpenStudy (danjs):

It looks like they rounded the answer, try moving the decimal one place to the left, then rounding to 1.0

OpenStudy (anonymous):

Alright, so the answer would be B.?

OpenStudy (danjs):

right 9.95 x 10^-19 = 0.995 x 10^-18 is about 1.0 x 10^-18

OpenStudy (anonymous):

Ok, cool. Thanks alot for the help. I really appreciate it.

OpenStudy (danjs):

welcome